locus of the centre of circles touching angle bisectors of the pair of lines ax2+bxy+ay2=0is..............???


    (answer--xy=0)

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear shivanshi

ax2+bxy+ay2=0

equation of the bisector of hese lines are

(x2 -y2) /(a-a) =xy/(b/2)

or

x2 -y2=0

y =x or y=-x

clearly these lines are symmetric about y axis and x axis

so if a circle touches both these two lines then its center must lie either on x axis or in y axis

so locus are  x=0

           and    y=0

 combine equation  is xy=0

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very

quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin